59395839495:

Given the expression: 5x10 − 80x2 Part A: Rewrite the expression by factoring out the greatest common factor. (4 points) Part B:Factor the entire expression completely. Show the steps of your work. (6 points)

59395839495:

pls give me the answer.

surjithayer:

\[5x^{10}-80x^2=5x^2(x^8-16)\] \[=5x^2[(x4)^2-4^2]\]\[=5x^2(x^4+4)(x^4-4)\] \[=5x^2(x^4+4)[(x^2)^2-2^2]\]\[=5x^2(x^4+4)(x^2+2)(x^2-2)\] \[=5x^2(x^4+4)(x^2+2)[x^2-(\sqrt{2})^2]\] \[=5x^2(x^4+4)(x^2+2)(x+\sqrt{2})(x-\sqrt{2})\]

LILBOY:

dude wht is this n goggle can help